The measure of an angle is 9 degrees more than twice its complement, find the measure of each angle.
Answer #1
x+(x+9)=90 2x+9=90 2x=90-9 2x=81 x=81/2 x=40.5 for the smaller angle. x+9-49.5 proof 40.5+49.5=90 90=90
answered by: bella's mom